# Comparative Analysis of $k$-essence and Quintessence Scalar Field Models: A Data Analysis Approach

### v1

@@ III.1 Dynamical system stability of $k$-essence field
The phase space of the system has been plotted for $\lambda=0.2$ in Fig. [3], where the gray region shows the constraint on field energy density, $(0\leq\Omega_{\phi}\leq 1)$. The green region shows the accelerating characteristics, $(-1\leq\omega_{\rm eff}<-1/3)$, when field energy density dominates $0.5<\Omega_{\phi}\leq 1$, and the phantom region $(-1.5<\omega_{\rm eff}<-1)$ is shown in red color. From the phase space, it becomes clear that the trajectories originate from $P_{1}$, which shows radiation-type characteristics, get attracted towards both $P_{3}$ and $P_{4}$. Those trajectories whose slopes are positive, $x>0$ are attracted towards $P_{4}$, while the trajectories with negative slopes, $x<0$, get attracted towards $P_{3}$. Depending on the choice of $\lambda$, either $P_{3}$ or $P_{4}$ shows accelerating characteristics. There are a few trajectories which are not constrained by the Hubble equation and are not attracted towards $P_{3,4}$; they get attracted to $P_{5}$, which shows an attractor matter regime. Note that point $P_{2}$ cannot be considered as a physically viable point, since $\Omega_{m}$ becomes negative for $\lambda^{2}<3$. In Fig. [4], we numerically evolved the system by varying the initial conditions and model parameters in the range of $0.5\leq x_{0}\leq 1.12$, $0.1\leq\Omega_{m0}\leq 0.32$, $0.001\leq\lambda\leq 0.4$ by generating random numbers. Here, we have varied $\Omega_{m0}$ instead of $y_{0}$, since one can get $y_{0}$ using the Hubble constraint equation as,
@@
Here, we can see that the initial conditions with $x_{0}>0$ remain positive; hence, the system stabilizes at point $P_{4}$ and generates an accelerating solution. We see that although during the late-time epoch, the field energy density dominates, but in the past epoch, the curves corresponding to the Hubble evolution show deviation from the data points. However, there is no substantial difference that can be accounted for in the distance modulus plot. In the subsequent sec. IV, we will conduct statistical analysis to obtain the best fit values of the dynamical variables and model parameters.
@@
Figure 4: The evolution of cosmological parameters of the k-essence field for $H_{0}=66\ \rm Km\ s^{-1}Mpc^{-1}$.
Model Parameters priors OHD OHD+Pantheon
$\Lambda$CDM $H_{0}$ $[40,100]$ $70.6\pm 1.5$ $66.8\pm 1.0$
$\Omega_{\Lambda 0}$ $[0.5,0.99]$ $0.742^{+0.022}_{-0.019}$ $0.675\pm 0.015$
$M$ $[-20,-19]$ $-19.454\pm{0.029}$
Quintessence $H_{0}$ $[40,100]$ $67^{+3.7}_{-2.9}$ $66.49\pm 0.98$
$x_{0}$ $[0,0.45]$ $0.266^{+0.17}_{-0.086}$ $0.311^{+0.036}_{-0.047}$
$y_{0}$ $[0.7,0.92]$ $0.810^{+0.059}_{-0.031}$ $0.804^{0.012}_{-0.018}$
$\lambda$ $[0,1.86]$ $1.04^{0.61}_{-0.41}$ $1.20^{0.11}_{-0.15}$
$M$ $[-20,-19]$ $-19.441\pm 0.029$
$k$-essence $H_{0}$ $[40,100]$ $66.7^{+3.1}_{-3.7}$ $66.41\pm 0.96$
$x_{0}$ $[0.7,1.4]$ $1.122\pm 0.086$ $1.118\pm 0.023$
$\Omega_{m0}$ $[0,0.9]$ $0.191^{+0.11}_{-0.060}$ $0.183^{0.12}_{-0.054}$
$\lambda$ $[-0.2,19]$ $0.30\pm 0.23$ $0.313^{0.065}_{-0.055}$
$M$ $[-20,-19]$ $-19.445\pm 0.028$
Table 3: The 68% confidence of the model parameters.
Model Data AIC BIC
$\Lambda$CDM OHD $27.63$ $31.15$
Pantheon+OHD $1794.26$ $1810.66$
Quintessence OHD $30.86$ $37.91$
Pantheon+OHD $1780.78$ $1808.10$
$k$-essence OHD $30.72$ $37.76$
Pantheon+OHD $1780.73$ $1808.05$
Table 4: The statistical quantities corresponding to the data sets.
Figure 5: The Quintessence field is fitted with OHD data.
Figure 6: The Quintessence field is fitted with OHD+Pantheon data.
Figure 7: The $k$-essence field is fitted with OHD data.
Figure 8: The $k$-essence field is fitted with OHD+Pantheon data.
IV Data Analysis
@@
Our subsequent objective involves performing data analysis to constrain the parameter space of the model. In order to achieve this goal, we need to derive an analytical expression for the Hubble parameter as a function of redshift. While this process is straightforward for the $\Lambda$CDM model, it poses challenges for quintessence and $k$-essence models due to the inability to obtain an analytical solution to the Friedmann equation. Therefore, we opted for a dynamical systems approach, solving the coupled differential equations outlined in Eqs. (14) and (32) to capture the evolution of the Hubble parameter. We utilize the most recent compilation of Type Ia supernovae (SNe Ia) from the pantheon+ dataset [54] along with Observational Hubble data (OHD) [55, 56, 57] to obtain the parameters of the model. We adopt the Bayesian parameter inference which relies on the Bayes theorem, offering a gratifying description to compute the posterior probability of a particular parameter of interest given the data (D) and model (M). Mathematically, it can be expressed as [58, 59],
@@
Here, the $\chi^{2}$ for the observational Hubble data is computed using the expression
$\chi^{2}_{\text{OHD}}=\sum_{k}\left[\frac{H_{k}-H_{k}(\theta)}{\sigma_{k}}\right]^{2},$ (37)
where $H_{k}$ is the Hubble parameter obtained from the data for the redshift $z_{k}$ and $H_{k}(\theta)$ is the corresponding Hubble parameter computed for the model. The OHD data consists of 43 Hubble parameter values obtained in the redshift span $2.360.07<z<$ 2.36. Among these, 31 data points are obtained from the differential age technique, which are model independent. The remaining 12 data points are derived from the Baryon acoustic oscillation measurement assuming the standard cosmology [56, 53]. The pantheon+ compilation include 1701 SNe Ia light curves of 1550 supernovae observed in the redshift span $0\leq z\leq 2.3$ [54]. The observable in case of SNe Ia is the apparent magnitude which is computed using the expression,
$m(z)=5\,log_{10}\left[\frac{d_{L}(z)}{Mpc}\right]+M+25,$ (38)
where $M$ represents the absolute magnitude of the SNe Ia. The luminosity distance $d_{L}$ in a flat FLRW universe can be calculated as [61],
$d_{L}(z)=c(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})},$ (39)
$\chi^{2}_{\text{SNe Ia}}=\Delta\vec{D}^{T}C^{-1}_{stat+syst}\Delta\vec{D},$ (40)
where $C_{stat+syst}=C_{stat}+C_{syst}$ is the combined covariance matrix of systematic and statistical uncertainties and $\vec{D}$ is the vector of 1701 SNe Ia distance modulus calculated as
$\Delta\vec{D}=\mu(z_{i})-\mu_{th}(z_{i},H_{0},\Omega_{m_{0}},M,y_{0}).$ (41)
Here, $\mu(z_{i})=m(z_{i})-M$ is the distance modulus of SNe Ia. The best fit values of the model parameters are those that minimize the total $\chi^{2}$,
$\chi^{2}_{\text{total}}=\chi^{2}_{\text{OHD}}+\chi^{2}_{\text{SNe Ia}}$ (42)
In order to obtain the 1D and 2D posterior probability distribution of the model parameters, we marginalize over the parameters except the parameters under consideration. The information criteria such as Akaike Information Criterion (AIC) [62] and Bayesian Information Criterion (BIC) [63] are used to compare the evidence of Quintessence or $k$-essence models over the $\Lambda$CDM model. The AIC and BIC are defined as [58],

### v2

@@ II.1 Stability of the system in the presence of Inverse square potential
Figure 3: The evolution of cosmological parameters of the k-essence field with power law potential for $H_{0}=70\ \rm Km\ s^{-1}Mpc^{-1}$.
The phase space of the system has been plotted for $\lambda=0.2$ in Fig. [2], where the gray region shows the constraint on field energy density, $(0\leq\Omega_{\phi}\leq 1)$. The green region shows the accelerating characteristics, $(-1\leq\omega_{\rm eff}<-1/3)$, when field energy density dominates $0.5<\Omega_{\phi}\leq 1$, and the phantom region $(-1.5<\omega_{\rm eff}<-1)$ is shown in red color. From the phase space, it becomes clear that the trajectories originating from $P_{1}$ (where near $P_{1}$ we have radiation-type characteristics) get attracted towards both $P_{3}$ and $P_{4}$. Those trajectories whose slopes are positive, $x>0$ are attracted towards $P_{4}$, while the trajectories with negative slopes, $x<0$, get attracted towards $P_{3}$. Depending on the choice of $\lambda$, either $P_{3}$ or $P_{4}$ shows accelerating characteristics. There are a few trajectories which are not constrained by the Hubble equation and are not attracted towards $P_{3,4}$; they get attracted to $P_{5}$, which shows an attractor matter regime. Note that point $P_{2}$ cannot be considered as a physically viable point, since $\Omega_{m}$ becomes negative for $\lambda^{2}<3$. The trajectory that lies outside the constrained region shown in blue cannot be considered physically viable, as it violates the essential constraints imposed on the system by the Hubble equation, Eq. (12). In Fig. [3], we numerically evolved the system by varying the initial conditions and model parameters in the range of
$0.8\leq x_{0}\leq 1.12,\quad 0.1\leq\Omega_{m0}\leq 0.31,\quad 0.01\leq\lambda\leq 0.3,$ (18)
by generating random numbers. Here, the subscript zero denotes the initial conditions at $N=\tilde{z}=0$. We have also plotted the evolution of the Hubble parameter $(H)$ in km/s/Mpc and distance modulus, $\mu$, against redshift, $\tilde{z}=1/(1+a)$, where $\mu$ is defined as
$\mu=5\log_{10}(d_{L})+25,$ (19)
where the luminosity distance is given by:
[…]
for the parameter ranges specified above. Here, $c$ represents the speed of light expressed in km/s. We compare it with the corresponding observational data including 43 OHD and 1701 Pantheon+SH0ES data sets [55, 56]. Note that, here, we vary $\Omega_{m0}$ instead of $y_{0}$, since one can get $y_{0}$ using the Hubble constraint equation as,
@@
In the figure, we can see that the initial conditions with $x_{0}>0$ produces values of $x$ that remain positive; hence, the system stabilizes at point $P_{4}$ and generates an accelerating solution. We see that although during the late-time epoch, the field energy density dominates, but in the past epoch, the curves corresponding to the Hubble evolution show deviation from the data points. However, there is no substantial difference that can be accounted for in the distance modulus plot. In the subsequent sec. III, we will conduct statistical analysis to obtain the best fit values of the dynamical variables and model parameters.
II.2 Stability of the system in the presence of exponential potential
In this section, we present the dynamical aspects of the $k$-essence field in the presence of an exponential type potential. The form of the potential is given by:
$V(\phi)=V_{0}e^{\beta H_{0}\phi}\,.$ (22)
Here, $V_{0}$ and $H_{0}$ are dimensional constants with mass dimensions $[M]^{4}$ and $[M]$, respectively, whereas $\beta$ is a dimensionless constant. To study the dynamics corresponding to the present case, we define a new set of dimensionless dynamical variables. These dynamical variables are different from the former case, although in some cases the notations are the same. The variables are:
$x=\dot{\phi},\quad y=\phi H_{0},\quad z=\frac{H-H_{0}}{H+H_{0}}\,,$ (23)
These are the primary variables required to close the system. Here, $H_{0}$ denotes the present value of the Hubble parameter. In the previous case, where the potential was considered of the inverse square type, the derivative of the potential, denoted as $\lambda$, turned out to be constant. As a result, only two variables were required to close the system. In contrast, the exponential type potential does not yield a constant $\lambda$ and thus requires more variables to close the system. Therefore, we have chosen the variables such that the dynamics of the system can be analyzed with a minimal number of dimensionless dynamical variables. It should be noted that the variables $(x,y)$ can take any value, as the range is given by:
$-\infty<x<\infty,\quad-\infty<y<\infty,$ (24)
whereas we have compacted the variable $z$ in the range given by:
$z=\begin{cases}-1&H\ll H_{0},\\ 0&H=H_{0},\\ 1&H\to+\infty\ .\end{cases}$ (25)
The secondary variables, for instance, the fractional densities of the field and fluid, become:
$\Omega_{\phi}:=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{3x^{4}}{4}-\frac{x^{2}}{2}\right),\quad\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}}\,.$ (26)
Here, $\mathcal{M}_{0}\equiv\frac{\kappa^{2}V_{0}}{H_{0}^{2}}$ is a dimensional constant. The Hubble equation puts a constraint on $\Omega_{m}$, which must be in the range given by:
$\Omega_{m}=1-\Omega_{\phi},\quad 0\leq\Omega_{m}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1.$ (27)
Because of the Friedmann constraint relation, the variables $(x,y)$ cannot take any values. Solutions that do not satisfy the above relation cannot be considered physically viable. Similarly, the effective equation of state (EoS) can be expressed in terms of the dynamical variable as:
$\omega_{\rm eff}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right).$ (28)
The autonomous system of equations can be constructed in the predefined variables as:
$x^{\prime}$ $=$ $\frac{-3x^{3}+\frac{\beta\left(\frac{x^{2}}{2}-\frac{3x^{4}}{4}\right)(1-z)}{z+1}+3x}{3x^{2}-1},$ (29)
$y^{\prime}$ $=$ $\frac{x(1-z)}{z+1},$ (30)
$z^{\prime}$ $=$ $-\frac{3((1-z)(z+1))\left(\frac{1}{3}\mathcal{M}_{0}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right)\left(\frac{1-z}{z+1}\right)^{2}\exp(\beta y)+1\right)}{4}\,.$ (31)
Here, $()^{\prime}\equiv\frac{d()}{Hdt}$, where $Hdt=dN$. Since the minimum number of dynamical variables is three, the corresponding phase space becomes 3D. Although the system of equations is consistent at $z=1$, i.e., $H\gg H_{0}$, one may notice the pathology of the dynamical system at $z=-1$, i.e., $H\ll H_{0}$. To study the dynamics corresponding to this coordinate, we use the following standard trick usually used in dynamical systems analysis by redefining the time variable as [48]:
$dN\to(1+z)d\tilde{N}.$ (32)
With this redefinition, the autonomous system of equations becomes non-divergent. The redefinition does not alter the dynamics of the system as long as the system is far from this critical point. However, in the data analysis section, to constrain the parameters, we shall evolve the system against the original time variable $dN=d\log(1/(1+\tilde{z}))$, where $\tilde{z}$ is the redshift. The experimental data available to us is in the redshift range $\tilde{z}\in[0.001,2.5]$, whereas $(\tilde{z}\to-1)$ shows the future state of the universe.
In the present case a conventional critical point analysis does not shed much light on the accelerating expansion phase of the universe and consequently we proceed with direct numerical evolution of the system using the autonomous equations. We take a different route to study the dynamics of the system. We skip the critical point analysis and numerically evolve the system by varying the parameters and initial conditions to obtain the dynamics of the system. We vary the appropriate variables within this range by generating random numbers as given by:
$0.21.01<x_{0}<1.25,\quad 0.3<\Omega_{m0}<0.31,\quad z_{0}=0,\quad 5<\mathcal{M}_{0}<9,\quad-0.2<\beta<$ 0.2 . (33)
Here, the subscript 00 indicates the value at the present epoch, for instance $x_{0}\equiv x(\tilde{N}=0)$. We have fixed $z_{0}=0$ due to its definition in Eq. (23). In this approach, we choose to vary $\Omega_{m0}$ instead of $y_{0}$ since, by following the Hubble constraint equation $\Omega_{m0}=1-\Omega_{\phi 0}$, one obtains the initial conditions for $y_{0}$. Note that this technique has also been adopted in the data analysis section III.
@@ II.2 Stability of the system in the presence of exponential potential
Figure 4: The evolution of cosmological parameters of the k-essence field for an exponential potential.
Figure 5: The evolution of cosmological observable $H$ in km/s/Mpc and distance modulus $\mu$ against redshift $\tilde{z}$ for the exponential potential. Here we considered the value of $H_{0}=70$ km/s/Mpc.
The summary of the evolution of cosmological parameters is illustrated in Fig. [4]. From the evolution of fractional densities of the field and fluid, it can be concluded that for a certain number of e-folds, $-3<\tilde{N}<-0.5$, the fluid energy density dominates over the field energy density, and the corresponding effective EoS $(\omega_{\rm eff})$ vanishes. This indicates that the matter phase becomes an attractor phase for the $k$-essence field with an exponential potential over a wide parameter range. In the far past epoch, we see that $\omega_{\rm eff}$ saturates to $1/3$ and the corresponding field energy density dominates. This suggests that the $k$-essence field with an exponential potential can produce a similar effect as a radiation fluid. As the system evolves towards the current epoch $\tilde{N}=0$, the field fractional density starts dominating, and the corresponding effective EoS decreases towards negative values, exhibiting a tracking behavior. The equation of state of the field becomes smaller than the EoS of the background fluid, causing the system to accelerate, which exhibits dark energy behavior. The dynamical variable $x$ corresponding to this phase nearly saturates to $1$. Throughout the evolution, the other dynamical variables do not diverge but take different values as the model parameters change.
As the system enters the far future $\tilde{N}>0$, the effective EoS $\omega_{\rm eff}$ for some combination of parameters even crosses $-1$, exhibiting phantom behavior, and the corresponding dynamical variable $z$ approaches $-1$. However, an additional interesting behavior can be observed when some curves of $\omega_{\rm eff}$ move towards positive values, hinting at a change of phase from an accelerating to a decelerating epoch. This scenario is also confirmed from the best fit obtained in Fig. [6b] for various combination of data sets. At this stage, the field density remains dominant over the fluid density.
Hence, from this analysis, it can be concluded that the $k$-essence field with an exponential potential can generate different phases of the universe, namely:
[…]
Additionally, we have plotted the Hubble parameter $H$ and distance modulus $\mu$ in Fig. [5], with the redshift $\tilde{z}$ assuming $H_{0}=70$ km/s/Mpc. It can be seen that the current model produces curves close to the observational data points. However, the best fit values of the parameters can only be obtained by accounting for MCMC with different data sets, which will be presented in the next section.
III Data Analysis
@@
Our subsequent objective involves performing data analysis to constrain the parameter space of the model. In order to achieve this goal, we need to derive an analytical expression for the Hubble parameter as a function of redshift. While this process is straightforward for the $\Lambda$CDM model, it poses challenges for $k$-essence models due to the inability to obtain an analytical solution to the Friedmann equation. Therefore, we opted for a dynamical systems approach, solving the coupled differential equations outlined in Eq. (15) to capture the evolution of the Hubble parameter. Here, we constraint the model parameters, utilizing the following data sets:
  • The Hubble datasets: This dataset contains $31$ points of Hubble parameter which are derived from the differential age method [57]. We refer to this dataset as CC. The chisquare ($\chi^{2}$) for this dataset is computed using the expression,
    $\chi^{2}=\sum_{i=1}^{31}\left(\frac{H_{i\ \rm obs}-H_{i\ \rm Model}}{\sigma_{i}}\right)^{2},$ (34)
    where $H_{i\ \rm obs}$ is the Hubble parameter observed at redshift $z_{i}$, $\sigma_{i}$ is the error in the measurement and $H_{i\ \rm Model}$ is the corresponding model prediction.
  • Pantheon+SH0ES data: This data set contains most recent compilation of Type Ia supernovae (SNe Ia) from the Pantheon+SH0ES dataset [58, 59]. The pantheon+ compilation include 1701 SNe Ia light curves of 1550 supernovae observed in the redshift span $0\leq\tilde{z}\leq 2.3$ [56]. It is pertinent to remind the reader that in our notation the cosmological redshift is specified by $\tilde{z}$ and not simply $z$ as is the conventional practice. We refer to this data set as PP throughout the text. The observable in case of SNe Ia is the distance modulus ($\mu$) which can be computed theoretically using the expression,
    $\mu(\tilde{z})=5\,\log_{10}\left[\frac{d_{L}(\tilde{z})}{Mpc}\right]+25,$ (35)
    The luminosity distance $d_{L}$ in a flat FLRW universe can be calculated as [60],
    $d_{L}(\tilde{z})=c(1+\tilde{z})\int_{0}^{\tilde{z}}\frac{dz^{\prime}}{H(z^{\prime})},$ (36)
    $\chi^{2}_{\text{SNe Ia}}=\Delta\vec{D}^{T}C^{-1}_{\rm stat+syst}\Delta\vec{D},$ (37)
    where $C_{\rm stat+syst}=C_{\rm stat}+C_{\rm syst}$ is the combined covariance matrix of systematic and statistical uncertainties and $\vec{D}$ is the vector of SNe Ia distance modulus calculated as
    $\Delta\vec{D}=\mu(\tilde{z}_{i})-\mu_{th}(\tilde{z}_{i},H_{0},\Omega_{m_{0}},y_{0}).$ (38)
  • DES data: The DES data set include 1829 SNe Ia in the redshift range $1.130.01<\tilde{z}<$ 1.13 [61].
  • BAO data: The data set contains 8 data points from Sloan Digital Sky Survey (SDSS) [62, 63] called as BAO in the following discussion and $7$ data points from Dark Energy Spectroscopic Instrument (DESI) [64, 65, 66] called as DESBAO in the following discussion, which are tabulated in Tab. [2]. The $\chi^{2}$ is defined as:
    $\chi^{2}_{\rm BAO}=\sum_{i=1}^{N}\left(\frac{X_{i}}{\sigma_{i}}\right)^{2}$ (39)
    where $X$ is given by:
    $X_{i}=\text{Data}_{i}-\text{Model}_{i}\ ,$ (40)
    Here, the observables are for instance $D_{M}(z)$, which is the comoving angular-diameter distance defined as $D_{M}(\tilde{z})=\int_{0}^{\tilde{z}}c\frac{dz^{\prime}}{H(z^{\prime})}$, the Hubble distance is $D_{H}(\tilde{z})=c/H(\tilde{z})$ and the spherically-averaged distance $D_{V}(\tilde{z})$ is called dilation scale defined as:
    $D_{V}(\tilde{z})\equiv\left({\tilde{z}\ D_{M}^{2}(\tilde{z})D_{H}}\right)^{1/3}\ .$ (41)
    The parameter $r_{d}$ shows the distance traveled by sound waves between the end of inflation and the decoupling of baryons from photons after recombination,
[…]
    where $c_{s}$ denotes the sound speed and $\tilde{z}_{d}$ is the redshift of the drag epoch. $c_{s}$ depends on the baryon and photon density which can be calculated as:
    $c_{s}(\tilde{z})=\frac{c}{\sqrt{3(1+\frac{3\rho_{B}}{4\rho_{\gamma}})}}\,.$ (43)
    However, in this study, we treat $r_{d}$ as a free parameter which we have constrained using the present combination of data sets. Note that to calculate the $\chi^{2}_{\rm BAO}$, we used $\chi^{2}_{\rm tot}=\frac{1}{2}\left(\chi^{2}_{D_{M}/r_{d}}+\chi^{2}_{D_{H}/r_{d}}+\chi^{2}_{D_{V}/r_{d}}\right)$ [62].
    BAO measurement
    $\tilde{z}_{\rm eff}$ 0.15 0.38 0.51 0.70 0.85 1.48 2.33 2.33
    $D_{V}(\tilde{z})/r_{d}$ $4.47\pm 0.17$ $18.33^{+0.57}_{-0.62}$
    $D_{M}(\tilde{z})/r_{d}$ $10.23\pm 0.17$ $13.36\pm 0.21$ $17.86\pm 0.33$ $30.69\pm 0.80$ $37.6\pm 1.9$ $37.3\pm 1.7$
    $D_{H}(\tilde{z})/r_{d}$ $25.00\pm 0.76$ $22.33\pm 0.58$ $19.33\pm 0.53$ $13.26\pm 0.55$ $8.93\pm 0.28$ $9.08\pm 0.34$
    DESBAO measurements
    $\tilde{z}_{\rm eff}$ 0.295 0.510 0.706 0.930 1.317 1.491 2.330
    $D_{V}(\tilde{z})/r_{d}$ $7.93\pm 0.15$ $26.07\pm 0.67$
    $D_{M}(\tilde{z})/r_{d}$ $13.62\pm 0.25$ $16.85\pm 0.32$ $21.71\pm 0.28$ $27.79\pm 0.69$ $39.71\pm 0.94$
    $D_{H}(\tilde{z})/r_{d}$ $20.98\pm 0.61$ $20.08\pm 0.60$ $17.88\pm 0.35$ $13.82\pm 0.42$ $8.52\pm 0.17$
    Table 2: BAO data.
We use the Bayesian parameter inference which relies on the Bayes theorem, offering a gratifying description to compute the posterior probability of a particular parameter of interest given the data (D) and model (M). Mathematically, it can be expressed as [67, 68],
@@
The best fit values of the model parameters are those that minimize the total $\chi^{2}$. In order to obtain the 1D and 2D posterior probability distribution of the model parameters, we marginalize over the parameters except the parameters under consideration. Information criteria, such as the Akaike Information Criterion (AIC) [70] and the Bayesian Information Criterion (BIC) [71], are utilized to compare how well $k$-essence models fit the data in comparison to the $\Lambda$CDM model. The AIC and BIC are defined as [67],